1. Consider all samples of size 4 from this population without replacement: 6 8 10 12 13 What is the mean and standard deviation of the Sampling Distribution? 2. Given the population 1, 3, 4, 6, and 8. Suppose samples of size 3 with replacement were drawn from this population. What is the mean and standard deviation of the Sampling Distribution? 3. A random sample of n=100 measurements is obtained from a population of N=120, with
1
Expert's answer
2021-10-27T10:21:20-0400

1.

 We have population values "6,8,10,12,13," population size "N=5" and sample size "n=4." Thus, the number of possible samples which can be drawn without replacement is



"\dbinom{N}{n}=\dbinom{5}{4}=5""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 6,8,10,12 & 9 \\
 \hdashline
 2 & 6,8,10,13 & 9.25 \\
 \hdashline
 3 & 6,8,12,13 & 9.75 \\
 \hdashline
 4 & 6,10,12,13 & 10.25 \\
 \hdashline
 5 & 8,10,12,13 & 10.75 \\
 \hline
\end{array}"

The sampling distribution of the sample means.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
& \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 & 9 & 1 & 0.2 & 1.8 & 16.2 \\
 \hdashline
 & 9.25 & 1& 0.2 & 1.85 & 17.1125 \\
 \hdashline
 & 9.75 & 1& 0.2 & 1.95 & 19.0125 \\
 \hdashline
 & 10.25 & 1& 0.2 & 2.05 & 21.0125 \\
 \hdashline
 & 10.75 & 1 & 0.2 & 2.15 & 23.1125 \\
 \hdashline
 Total & & 5 & 1 & 9.8 & 96.45 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=9.8"




"\mu=\dfrac{6+8+10+12+13}{5}=9.8"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=\mu_{\bar{X}}=9.8=\mu"




"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=96.45-(9.8)^2=0.41"




"\sigma_{\bar{X}}=\sqrt{Var(\bar{X})}=\sqrt{0.41}\approx0.6403"

Verification:


"\sigma^2=\dfrac{1}{5}((6-9.8)^2+(8-9.8)^2+(10-9.8)^2"

"+(12-9.8)^2+(13-9.8)^2)=6.56"


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{6.56}{4}(\dfrac{5-4}{5-1})""=\dfrac{6.56}{16}=0.41, True"


2.

 We have population values "1,3,4,6,8," population size "N=5" and sample size "n=3." Thus, the number of possible samples which can be drawn with replacement is



"N^n=5^3=125"


The mean of the sampling distribution of the sample means is equal to the the mean of the population.


"\mu_{\bar{X}}=\mu=\dfrac{1+3+4+6+8}{5}=4.4"

"\sigma^2=\dfrac{1}{5}((1-4.4)^2+(3-4.4)^2+(4-4.4)^2"

"+(6-4.4)^2+(8-4.4)^2)=5.84"

"Var(\bar{X})=\sigma^2_{\bar{X}}=\dfrac{\sigma^2}{n}=\dfrac{5.84}{4}=1.46"

"\sigma_{\bar{X}}=\sqrt{Var(\bar{X})}=\sqrt{1.46}\approx1.2083"

3.

We have population size "N=120" and sample size "n=100." 

The mean of the sampling distribution of the sample means is equal to the the mean of the population.


"\mu_{\bar{X}}=\mu"

Variance of the sampling distribution of the sample means is


"Var(\bar{X})=\sigma^2_{\bar{X}}=\dfrac{\sigma^2}{n}"

Standard deviation of the sampling distribution of the sample means is


"\sigma_{\bar{X}}=\sqrt{Var(\bar{X})}=\dfrac{\sigma}{\sqrt{n}}"